A solution is prepared by dissolving a certain amount of solute in 500 g of water. The percentage by mass of a solute in a solution is 2.38. Calculate the mass of solute. 

A

10.42 g

B

11.45 g

C

12.19 g

D

14.22 g

 

Grade:10

4 Answers

ankit singh
3 years ago

Weight percentage=[Weight of solute+ weight of solventWeight of solute]×100 

Let the weight of the solute be 'x'.
2.38=x+500x×1002.38x+1190=100x1190=97.62xx=97.621190=12.190g
 Mass of solute =12.19 grams
ankit singh
3 years ago
Weight percentage=[ 
Weight of solute+ weight of solvent
Weight of solute
 ]×100 
Let the weight of the solute be 'x'.
2.38= 
x+500
x×100
 
⟹2.38x+1190=100x
⟹1190=97.62x
⟹x= 
97.62
1190
 =12.190g
∴ Mass of solute =12.19 grams
irshad hussain
2 years ago
correct option is
C

12.19 g

Weight percentage=[Weight of solute+ weight of solventWeight of solute]×100 

Let the weight of the solute be 'x'.
2.38=x+500x×1002.38x+1190=100x1190=97.62xx=97.621190=12.190g
 Mass of solute =12.19 grams
irshad hussain
2 years ago
orrect option is
C

12.19 g

Weight percentage=[Weight of solute+ weight of solventWeight of solute]×100 

Let the weight of the solute be 'x'.
2.38=x+500x×1002.38x+1190=100x1190=97.62xx=97.621190=12.190g
 Mass of solute =12.19 grams